Peter Sagan is in charge of maintaining hospital supplies at Champs Hospital. During the past year the mean weekly demand for a special type of tubing was 186 packages of this tubing with a standard deviation of 13 packages of tubing. The lead time for receiving this tubing from the supplier is 1.5 weeks. Peter would like to maintain a 95% service level and places an order for 750 packages every time an order is placed.

a) How much safety stock should be used for a 95% service level?

b) If the weekly demand is 186 and there is a 1.5 week lead time - what is the reorder point (95% service level)?

c) How much safety stock should be used for a 99% service level?

d) If the weekly demand is 186 and there is a 1.5 week lead time - what is the reorder point (99% service level)?

e) If the carrying cost per year is $0.50/unit/year - what is the additional cost associate with the 99% service level compared to 95% service level (i.e. cost of safety stock at 99% level - cost of safety stock at 95% service level)?
Answer #1

Reorder point = average weekly demand * lead time + safety stock

Where safety stock = Z-score * Standard deviation * sqrt( lead time)

Z-score for 95% service level = 1.65, Z-score for 99% level = 2.33

a) Safety stock = 13*1.65*sqrt(1.5) = 26.27

b) Reorder point 95% service level =186*1.5+26.27 = 305.27 =~305 units

c) Safety stock at 99% stock = 2.33*13*sqrt(1.5) = 37.09

d) Reorder point = 186*1.5+37.09 = 316.09 =~ 316 units

Note - Answered 1st 4 parts as per Q&A policy.